OpenStudy (anonymous):

write a quadratic equation whose roots are -1 and 2 and whose leading coefficient is 2

OpenStudy (nikato):

because the roots are -1 and 2 so (x+1)(x-2)=0 and multiply one of them by 2 because its a leading coefficient. lets say the first one 2(x+1)(x-2)=0 and distribute (2x+2)(x-2)=0 FOIL and combine like terms to find your equation